OpenStudy (anonymous):

Find the derivative of 3^(-x^2)

OpenStudy (anonymous):

\[\frac{ d}{ dx} \left[ 3-^{x ^{2}}\right]\]

OpenStudy (anonymous):

I thought the derivative of a^u is (a^u)(ln a)(du/dx) so I got 3^(-x^2) (ln3) (-2x) but the answer is -3^(-x^2)xln9

OpenStudy (anonymous):

http://www.derivative-calculator.net/ calculator for it ^

OpenStudy (anonymous):

Their method is too long and involves setting all of the input derivatives equal to imaginary y value and then making the problem in terms of e^logx whatever. Imo it's useless.

OpenStudy (anonymous):

Or needlessly complex and difficult to understand, rather.

jimthompson5910 (jim_thompson5910):

Let y = 3^(-x^2) Apply the natural log to both sides ln(y) = ln[ 3^(-x^2) ] ln(y) = -x^2*ln(3) Now derive both sides with respect to x and isolate dy/dx

OpenStudy (anonymous):

Is there a way to do it without implicit differentiation?

jimthompson5910 (jim_thompson5910):

you can use ` derivative of a^u is (a^u)(ln a)(du/dx) ` u = -x^2 du/dx = -2x a = 3 y = a^u dy/dx = (a^u)(ln a)(du/dx) so if y = 3^(-x^2) then dy/dx = (a^u)(ln a)(du/dx) dy/dx = 3^(-x^2)*ln(3)*(-2x) dy/dx = -3^(-x^2)*x*2ln(3) dy/dx = -3^(-x^2)*x*ln(3^2) dy/dx = -3^(-x^2)*x*ln(9)

OpenStudy (anonymous):

Oh. I forgot that you could apply the log power rule here. Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

Here is how to do it using implicit differentiation Let y = 3^(-x^2) ln(y) = ln[ 3^(-x^2) ] ln(y) = -x^2*ln(3) y'/y = -2x*ln(3) y' = y*(-2x*ln(3)) y' = -3^(-x^2)*2x*ln(3) y' = -3^(-x^2)*x*2ln(3) y' = -3^(-x^2)*x*ln(3^2) y' = -3^(-x^2)*x*ln(9) so we get the same answer